OpenStudy (anonymous):

Solve the linear inequality and express your answer in interval notation. -3|x-13|+33 is less than or equal to 3

OpenStudy (anonymous):

haha thanks

OpenStudy (anonymous):

I suck at those :(

OpenStudy (anonymous):

samee

OpenStudy (anonymous):

hold on a second

OpenStudy (anonymous):

\[-3|x-13| \le -30\] \[|x-13|\geq 30\] right?

OpenStudy (anonymous):

lemme check

OpenStudy (anonymous):

no

OpenStudy (jamesj):

@sat73 is on the case again and correct.

OpenStudy (jamesj):

You can see this the solution as derived is wrong. Take for example x=10 that lies in the purported answer set. Calculate -3|x-13|+33 = -3|3| + 33 = -9 + 33 = 24 which is not less than 3.

OpenStudy (anonymous):

So from |x−13|≥10, x ≥ 23 or x =< 3. In interval notation (−∞,3]∪[23,∞)

OpenStudy (anonymous):

Thank you James. But, Simon thanks so much the answer was correct! I greatly appreciate it.

OpenStudy (jamesj):

Give 'im a medal!

OpenStudy (anonymous):

Sure will!